OpenStudy (anonymous):

General solution to a set of equations with infinite solutions... (see comments for specific qu) :)

OpenStudy (anonymous):

The following equations have infinitely many solutions. −15 y − 12 z = 6 −2 x − 13 y − 11 z = 5 15 y + 12 z = −6 Give the right hand side of the vector form of the general solution, using a parameter such as s or t. (Any lowercase letter will do as a parameter, so long as it is not x, y or z.)

OpenStudy (anonymous):

An example we have been given: for the equations x = y + 1 y = z + 1 z = x − 2 one correct answer is [x, y, z] = [0, −1, −2] + t [1, 1, 1] and you would enter the right hand side of this equation in the space provided: [0, −1, −2] + t * [1, 1, 1] or, if you prefer: [t, t − 1, t − 2] **NB "in the space provided" refers to an online-submission system for answer checking**

OpenStudy (zarkon):

\[\left[\begin{matrix}0 & -15 & -12 & 6 \\ -2 & -13& -11 & 5 \\0 & 15 & 12 & -6\end{matrix}\right]\] will row reduce to...

OpenStudy (zarkon):

\[\left[\begin{matrix}1 & 0 & 3/10 & 1/10 \\ 0 & 1& 4/5 & -2/5 \\0 & 0 & 0 & 0\end{matrix}\right]\]

OpenStudy (zarkon):

so we have one free variable

OpenStudy (anonymous):

yep, agreed :)

OpenStudy (zarkon):

we can use z=t for the free variable so the 2nd row gives us y+4/5t=-2/5 so y=-4/5t-2/5

OpenStudy (zarkon):

1st row gives us x+3/10t=1/10 so x=-3/10t+1/10

OpenStudy (zarkon):

so our solution is \[\left[\begin{matrix}x \\ y \\z \end{matrix}\right]=\left[\begin{matrix}3/10 \\ -4/5 \\1 \end{matrix}\right]t+\left[\begin{matrix}1/10 \\ -2/5 \\0 \end{matrix}\right]\]

OpenStudy (anonymous):

You are a legend! thanks so much!!

OpenStudy (zarkon):

np

OpenStudy (zarkon):

\[\left[\begin{matrix}x \\ y \\z \end{matrix}\right]=\left[\begin{matrix}-3/10 \\ -4/5 \\1 \end{matrix}\right]t+\left[\begin{matrix}1/10 \\ -2/5 \\0 \end{matrix}\right]\] left of a negative sign on the 3/10 (now fixed -3/10)

OpenStudy (anonymous):

haha yeah i was just scratching my head about that, too :P cheers!